XUV LASER PHOTOIONIZATION SPECTROSCOPY AT A RESOLUTION OF 0.008 CM^{−1}

\underline{U. HOLLENSTEIN}, G. C. L. PETRAGLIO, H. PALM, R. SEILER and F. MERKT, Laboratorium für Physikalische Chemie, Eidgenössische Technische Hochschule, ETH-Zentrum, CH-8092 Zürich, Switzerland.

A new broadly tunable (7-17 eV) narrow bandwidth (0.008 cm^{−1}) VUV laser system is presented. The system has been used to record photoionization spectra of small molecules (N_{2}, CO) in the vicinity of their ionization threshold. At the high resolution detailed information can be extracted on the spectroscopic and dynamic properties of electronically excited states, in particular Rydberg states.